distance between two points in a medium is 2m optical path corresponding to this distance is 3m. find out the velocity of light in that medium !

Grade:12th pass

1 Answers

Khimraj
5 years ago
optical path length = (geometric length)*(refractive index of medium)
3 = 2*c/v
v = (2/3)*3*108 m/s
so v = 2*108 m/s
Hope it clears. If you like answer then please approve it.